# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

### v1

@@ 1. Introduction
2. Related Works
@@
Initial resources for sentiment classification in English language comprises of product reviews (Pang et al., 2002; Maas et al., 2011), and the relevance of sentiment classification of product reviews will probably never diminish. However, extensive use of social media platforms has drawn attention to understanding sentiment polarity of short social media posts (Socher et al., 2013; Dong et al., 2014; Wang et al., 2010; Rosenthal et al., 2017; Orbach et al., 2021). Casual nature of such texts makes them extremely noisy containing grammatical errors, typos, an abundance of non-standard text, slang and code-mixing. This posed new challenges in curating benchmark datasets and addressing the problem itself. Mostly, sentiment classification has been considered as a binary (Socher et al., 2013; Orbach et al., 2021; Pang et al., 2002; Maas et al., 2011), 3-class (Dong et al., 2014), or 5-class (Wang et al., 2010; Rosenthal et al., 2017) problem. While most of the popular datasets sourced the data from one particular social media platform such as Twitter (Dong et al., 2014), IMDB (Maas et al., 2011), tripAdvisor (Wang et al., 2010), etc., datasets like (Orbach et al., 2021) considered diversifying sources over multiple platforms.
@@
The lack of high-quality datasets for sentiment classification in low-resource Indic languages has significantly limited research in this area. The most commonly cited corpora include Hindi (Akhtar et al., 2016), Bangla (Islam et al., 2023), Marathi (Kulkarni et al., 2021), multiple languages that include Telugu (Doddapaneni et al., 2023) and code-mixed texts with English (Patwa et al., 2020; Chakravarthi et al., 2020a, b). (Mukku and Mamidi, 2017) and (Marreddy et al., 2022) are the only popular Telugu sentiment corpus. Among the aforementioned Indic corpora, only Islam et al. (Islam et al., 2023) follow a well-defined annotation protocol, described in detail, and have curated a large-scale dataset from Bangla texts across diverse domains, collected from multiple social media platforms. The remaining studies did not diversify their data sources. A few of them have small sample sizes (Mukku and Mamidi, 2017; Chakravarthi et al., 2020a), some do not discuss the annotation protocol in detail (Marreddy et al., 2022; Kulkarni et al., 2021; Akhtar et al., 2016), and others suffer from weak data and annotations (Doddapaneni et al., 2023; Kulkarni et al., 2021; Patwa et al., 2020; Chakravarthi et al., 2020a; Mukku and Mamidi, 2017). None of these works including Islam et al. (Islam et al., 2023) extend their benchmarks to address explainability and fairness considerations. To the best of our knowledge, ours is the first large-scale Telugu sentiment corpus—and possibly the second among Indic languages after Islam et al. (Islam et al., 2023)—comprising texts from diverse domains, sourced from multiple social media platforms, and developed using a well-defined and well-documented annotation protocol. Furthermore, it is the first Indic sentiment corpus to incorporate considerations of explainability and fairness.
Explainability
@@
Several of-the-shelf explainers have gained popularity for model auditing which are post-hoc in nature (Bhatt et al., 2020). These explainers weight input tokens based on their contribution to the prediction. Since they are based on crude heuristics, these explainers are inherently imperfect (Rudin, 2019). Plausibility and Faithfulness are the two primary categories of evaluation measures for these explainers (Lyu et al., 2024). A salient feature of our dataset is the inclusion of human-annotated rationales, which serve as ground truth for plausibility evaluation. While this has been explored to some extent in English (Yeruva et al., 2020), our work is the first to address it in the context of an Indic language.
Fairness
@@
Fairness in sentiment classification has been evaluated in English for various social biases, including gender bias (Thelwall, 2018; Kiritchenko and Mohammad, 2018), age-related bias (D´ıaz et al., 2018), and racial bias (Kiritchenko and Mohammad, 2018) among others. Attempts to evaluate fairness in Indic sentiment classification are rare, with only a few recent studies—such as those in Bangla (Das et al., 2023, 2024)—curating dataset for addressing this issue. Our work is the first attempt to provide corpus and subsequently evaluate fairness in Telugu sentiment classification, and a rare effort for the same across Indic languages.
3. Data Collection & Preprocessing
To create a diverse corpus, we collected data from multiple sources: YouTube comments, comments from various Telugu blogs, news websites (Ennadu, Andhra Jyothi, Way2News (Telugu), and Sakshi) , and Facebook, Instagram comments.
For YouTube, we curated a set of standard topics (e.g. Politics, Sports, etc.) and search phrases per topic (details provided in Appendix F). Using the YouTube API library (version 3), we retrieved up to 50 top videos for each search phrase. We then extracted all comments that included at least one Telugu word or phrase using Langdetect at 90% confidence threshold. The whole extraction process took place between September 10th and 20th, 2024.
For news websites, using BeautifulSoup (version 4), we scraped article headlines from the aforesaid news websites over one week, spanning 25th September to 2nd October. Note, this refers to the scraping period, not the publication dates of the articles.
Lastly, we collected social media comments using the Instant Data Scraper web extension and with a few generic hashtags (e.g. #telugumemes, #teluguvideos etc) to crawl content from Facebook & Instagram. We scarped between the period of  10th Sept to 12th Sept. We adhered to the default limits and parameters while using every API, scraper, and crawler. We refer the readers to Appendix A for a quantitative overview of the data collection from various sources.

### v2

@@ 1 Introduction
2 Related Works
Sentiment classification has been extensively studied in English, beginning with product-review datasets Pang et al. (2002); Maas et al. (2011) and later extending to large-scale social media benchmarks that address informal, noisy text and multi-class sentiment settings Socher et al. (2013); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021). In contrast, sentiment analysis for Indic languages remains severely under-resourced. Existing Indic datasets span a limited number of languages and often suffer from small-scale, narrow domain coverage, weak or undocumented annotation protocols, or reliance on translated rather than native text Akhtar et al. (2016); Chakravarthi et al. (2020a); Patwa et al. (2020); Doddapaneni et al. (2023). Telugu, despite its large speaker base, is particularly underrepresented, with only a small number of sentiment corpora available Mukku and Mamidi (2017); Marreddy et al. (2022).
More broadly, existing Indic sentiment benchmarks do not support systematic investigation of human-centered supervision, explainability, or fairness. In English, text classification benchmarks that include human annotated rationales—including those for sentiment analysis—have primarily been used to improve model performance or to evaluate the plausibility of post-hoc explanation methods Herrewijnen et al. (2024). More recently, Explanation-Guided Learning (EGL) Gao et al. (2024) has emerged as a research direction that incorporates human-annotated explanations directly into training and studies their effects on model behavior, explanation quality, and fairness, accountability, and transparency (FAccT) properties. However, to the best of our knowledge, no prior work has systematically explored EGL in Indic languages, or more broadly in low-resource language settings. Our work addresses this gap by introducing a large-scale, human-centered Telugu sentiment resource explicitly designed to support supervision beyond labels, explainability analysis, and fairness evaluation.
The paradigm of human alignment is multi-facade and broadly refers to techniques for steering model behavior toward human expectations Christian (2020). Prior human-centered supervision paradigms such as instruction tuning Wei et al. (2022) and preference learning Ouyang et al. (2022) have predominantly been studied in the context of large language models performing text generation, where alignment is achieved through output-level or interaction-level feedback. In contrast, our work focuses on discriminative sentiment classification with pre-trained encoder-based models, incorporating annotator-provided rationales alongside sentiment labels as input-grounded supervision. This setting enables a systematic examination of how human-centered supervision influences classification behavior, the alignment of model explanations with human reasoning, and broader properties such as bias and fairness in a low-resource language.
A more detailed discussion of related benchmarks, explainability, and fairness literature is provided in Appendix B.
3 Data Collection & Preprocessing
To create a diverse corpus, we collected data from multiple sources: YouTube comments, comments from various Telugu blogs, news websites (Ennadu, Andhra Jyothi, Way2News (Telugu), and Sakshi).
For YouTube, we curated a set of standard topics (e.g. Politics, Sports, etc.) and search phrases per topic (details provided in Appendix M). Using the YouTube API library (version 3), we retrieved up to 50 top videos for each search phrase. We then extracted all comments that included at least one Telugu word or phrase using Langdetect at 90% confidence threshold. The whole extraction process took place between September 10th and 20th, 2024.
For news websites, using BeautifulSoup (version 4), we scraped article headlines from the aforesaid news websites over one week, spanning 25th September to 2nd October. Note, this refers to the scraping period, not the publication dates of the articles [^3] Along with this, we previously scraped Facebook and Instagram comments; however, due to ambiguities in Meta’s Terms of Service, these were ultimately excluded from the final version of TeSent..
The raw corpus was subsequently subjected to a comprehensive preprocessing pipeline involving language filtering, deduplication, and anonymization. For clarity and reproducibility, the complete preprocessing procedure is described in Appendix C. After preprocessing, from the annotated corpus of 23,986 sentences, 21,743 were found to be valid (see Appendix E). In addition, out of 607 sentences selected for testing, 482 were valid. This brought the total number of valid annotated sentences to 22,225.
4 Annotation Process
4.1 Overview
The annotation process was designed to capture both sentiment polarity and fine-grained human rationales in Telugu text. A pool of trained native Telugu speakers annotated each sentence using a structured, two-layer framework that balances annotation quality with cognitive efficiency. Details regarding annotator recruitment and our custom annotation software for doing it are provided in Appendix D and Appendix E, respectively.
4.2 Annotation Framework
In contrast to prior efforts, we adopt a distinct methodological framework for annotation. To support this, we also developed custom in-house annotation software tailored to our specific design principles, as detailed in Appendix E. Each annotator annotates every sentence following two layers of annotation. First, annotators are instructed to assign a primary sentiment label. Following this, they are asked to highlight specific words or phrases in the sentence that serve as human rationales for their chosen label. To capture subtler sentiment dynamics, annotators are then given the option to provide a secondary sentiment label if they believe one reasonably coexists with the primary annotation. However, for brevity and to reduce cognitive load, we do not collect rationales for this secondary label.
4.3 Finalizing the Label
To finalize the sentiment label for each sentence, we begin by discarding annotations where the primary and secondary sentiments are direct contradictions—specifically, cases where the primary label is positive, and the secondary is negative, or vice versa (we denote this as ‘<+,-> pair’). Such contradictions may indicate either annotator inconsistency or that the sentence was noisy or ambiguous to them, and including these could compromise label quality. Following this filtering step, we apply a majority voting scheme over the primary sentiment labels provided by all annotators. If a clear majority emerges from that, we assign that label as the final one. In cases where there is no majority, i.e., when annotators choose all three different labels (positive, negative, and neutral), we then incorporate the available secondary labels and re-run the majority vote. However, in such cases, we discovered that although secondary sentiments help resolve the no majority problem, the final label still could be confusing. Since such cases were rare in our dataset, we did not perform additional manual intervention and discarded the small split containing $\leq 3.2\%$ of the sentences from the final dataset.
Out of 22,225 total valid and annotated sentences, 19,755 had only primary annotations from all annotators, of which 19,087 had a clear majority winner and were retained. The remaining 2,470 sentences included at least one annotation with both a primary and secondary label; after removing 58 due to a technical glitch, 2,412 remained. From these, annotations with <+, -> pairs were removed where present, and majority voting over the priority labels yielded 2,115 sentences. This resulted in a corpus of 21,202 sentences. Finally, Telugu-speaking authors manually inspected the entire annotated dataset and, at their discretion, removed an additional 83 malformed or duplicate sentences, leading to a final total of 21,119. Overall, we retained over 95% of high-quality sentences with the final label from the set of valid sentences. Lastly, with input from a senior Telugu-speaking co-author and using the off-the-shelf IndicNER model, we performed named entity recognition and did not identify any offensive sentences to the best of our understanding.
We provide an overview of the dataset’s lexical and genre characteristics, including word-count statistics and genre distribution, in Appendix F.
4.4 Annotation Statistics and Interpretation
Our final dataset consists of 45.42% neutral, 26.84% positive, and 27.74% negative instances, annotated by a large pool of annotators with varying levels of contribution; additional details are provided in Appendix F. Inter-annotator agreement was measured using Krippendorff’s Alpha ($\alpha$) and Fleiss’s Kappa ($\kappa$) Gwet (2014). For the three-class sentiment setting, we obtain $\alpha=0.4706$ and $\kappa=0.4705$.
Interpreting agreement coefficients in multi-class settings is less straightforward than in binary tasks, and commonly used qualitative labels (e.g., “moderate” or “strong”) are primarily derived from binary or near-binary scenarios McHugh and others (2012). For this reason, we avoid assigning threshold-based qualitative interpretations to these values. Notably, the observed agreement is consistent with prior methodologically comparable studies; for example, Mathew et al. (2021) report a Krippendorff’s $\alpha$ of 0.46 for a three-class social-media annotation task with inherent ambiguity, which also incorporates human-annotated rationales alongside labels.
To provide a more interpretable view of annotator consistency, as suggested by prior work emphasizing complementary agreement analyses beyond summary coefficients McHugh and others (2012), we additionally report the cumulative distribution of pairwise agreement among annotators (Figure 1). The distribution shows that approximately 80% of annotators agree with the remaining annotators on more than 82% of instances, and nearly all annotators exceed a 75% agreement level, indicating that disagreement is concentrated in a relatively small subset of instances.
Finally, when collapsing the task to a binary sentiment setting by excluding neutral instances, agreement increases substantially to $\alpha=0.8928$, indicating strong consistency when sentiment polarity is unambiguous. Taken together, the binary agreement, the agreement distribution in the three-class setting, and consistency with prior work support the reliability of the annotation process for capturing nuanced sentiment distinctions in Telugu.
Figure 1: Distribution of Agreement Levels
5 Fairness Benchmarking
Sentiment classification systems can perpetuate various societal biases, including gender, racial, and identity-based biases. The presence of such bias often leads to disparities in classification accuracy or sentiment polarity across different demographic groups. Kiritchenko et al. Kiritchenko and Mohammad (2018) proposed the Equity Evaluation Corpus (EEC), an English dataset designed to evaluate age and racial bias across multiple NLP tasks related to sentiment and emotion. We extend and tune EEC accommodating niche morphological nuances in Telugu language to create TeEEC, Telugu Equity Evaluation Corpus, and propose a robust fairness evaluation framework for sentiment classification in Telugu, with a focus on gender and religion bias. Although our bias evaluation framework is specifically designed for sentiment classification, TeEEC can be applied to any NLP task involving sentiment or emotion.
5.1 TeEEC
We construct the Telugu Equity Evaluation Corpus (TeEEC) to evaluate demographic fairness in Telugu sentiment classification using a controlled, template-based counterfactual framework derived from Kiritchenko et al. Kiritchenko and Mohammad (2018). TeEEC is based on 11 sentence templates originally proposed in English, comprising seven templates with explicit emotion words and four sentiment-neutral templates without emotion words (e.g., “<person> feels <emotion word>” and “I saw <person> in the market”). In the former, the sentiment polarity is determined solely by the emotion word, while in the latter, any difference in prediction can be attributed only to the demographic reference. All templates are translated into Telugu while preserving their original structure and intent, and no new templates are introduced, ensuring consistency with prior work and avoiding additional sources of variation. For the emotion-bearing templates, we select emotion words from four affective categories-Anger, Fear, Joy, and Sadness-and choose ten commonly used Telugu emotion words from each category; the inserted emotion word determines the sentiment polarity of the sentence, with morphological realization chosen to ensure grammatical correctness. [^4] We consider two morphological realizations of emotion words—a noun form and a corresponding adjectival form—and select the form required by each template; a manual analysis shows that Templates 3, 4, and 7 (as numbered in Kiritchenko and Mohammad (2018)) license noun forms, while the remaining templates require adjectival forms.
TeEEC evaluates fairness with respect to gender $\{\text{male, female}\}$ and religion $\{\text{Hindu, Muslim, Christian}\}$, instantiated through the <person> placeholder using names or noun phrases associated with the corresponding group. For gender evaluation, we curate 20 commonly used Telugu names for each gender and 10 noun phrases per group, where the noun phrases exhibit a one-to-one correspondence across genders and do not encode religion; for religion evaluation, names are grouped by religious affiliation, and noun phrases are not used. We do not include caste, age, occupation, or socio-economic factors, although these are relevant in the Indian context, because they cannot be incorporated into this counterfactual framework without introducing ambiguity or severely limiting coverage. [^5] Caste is often not unambiguously identifiable from names and is frequently conflated with religion or region; for example, surnames such as Naidu or Reddy may correspond to different communities across regions or persist across religious conversion (e.g., John Reddy, Mary Kumari), making it infeasible to generate unambiguous counterfactual sentence pairs. Directly inserting caste labels would also yield very few usable counterfactuals; similar challenges arise for occupation and age groups. Moreover, since our goal is to evaluate bias specifically for sentiment classification, each template must include a sentiment identifier, which makes curating specialized templates for additional factors infeasible within the scope of this work. Manually curating or annotating sentences (including with LLMs) for bias evaluation is outside the scope of this study and is left for future work [^6] Hindu and Christian names were randomly sampled from the curator-maintained reference site Behind the Name, and Muslim names from the analogous resource QuranicNames. Within the best of our capacity, we verified whether sampled names appear in the respective religious texts using publicly available web sources; however, such verification was not possible for all names..
For gender bias evaluation, we generate two disjoint sentence sets corresponding to male and female references, each containing all possible combinations of demographic-specific <person> realizations and emotion words across all emotion-bearing templates, along with sentences derived from sentiment-neutral templates, resulting in a total of 39,760 sentences evenly split between the two gender groups; the religion bias dataset is constructed following the same procedure using names only. TeEEC thus provides, to the best of our knowledge, the first large-scale equity evaluation corpus for Telugu covering the two most prominent and identifiable societal dimensions—gender and religion—under a controlled and reproducible setup, and the complete dataset and accompanying Python scripts are publicly available at https://huggingface.co/datasets/DSL-13-SRMAP/Telugu-Equity-Evaluation-Corpus_TeEEC.
5.2 Evaluation Metrics
We evaluate bias in two distinct ways. First, we measure differences in predicted sentiment polarities across demographic groups using a counterfactual test Kiritchenko and Mohammad (2018); Goldfarb-tarrant et al. (2023), which we formalize as the Polarity Difference Score ($PDS$). Second, we assess fairness using equalized odds (EO) Hardt et al. (2016), computed as the absolute differences in true positive rates and false positive rates across groups, denoted as $TPR_{diff}$ and $FPR_{diff}$. Details of the counterfactual test, PDS, and EO are provided in Appendix G.
6 Explainability Benchmarking
Commonly used post-hoc local explainers can be broadly categorized into two types: perturbation-based and gradient-based explainers Arrieta et al. (2020). In this work, we consider two widely adopted perturbation-based, model-agnostic explainers: LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017), using the default selection of partition shap for SHAP.
Among the gradient-based methods, we include Gradient (Grad.) Simonyan et al. (2013), Integrated Gradients (Int. Grad.) Sundararajan et al. (2017), and their input-scaled variants: Gradient × Input (Grad. × Input) and Integrated Gradients × Input (Int. Grad. × Input).
To evaluate faithfulness, we compare our results using well-established erasure-based metrics: comprehensiveness (C($\uparrow$)), sufficiency (S($\downarrow$)) DeYoung et al. (2020), and correlation with Leave-One-Out (L($\uparrow$)) scores Jain and Wallace (2019).
For plausibility analysis, we followed DeYoung et al. DeYoung et al. (2020) and report token-level Intersection-Over-Union (I($\uparrow$)), token-level F1 scores (F($\uparrow$)), and Area Under the Precision-Recall Curve (A($\uparrow$)). For all measures, $\uparrow$’ signifies the higher the better and $\downarrow$’ signifies the lower the better.
We conduct all experiments using the ferret-xai library with its default (hyper)parameter settings Attanasio et al. (2023) and report the mean values for each measure.
7 Experimental Results
7.1 Model Training & Performance Analysis
[…]
Across models, training with rationale supervision generally improves performance compared to label-only supervision. Four of the five architectures (mBERT, IndicBERT, MuRIL, and Te-BERT) show higher F1 and AUROC, with the largest relative gains observed in MuRIL and Te-BERT; accuracy in these models also increases. IndicBERT shows smaller but consistent improvements, while XLM-R is the only model for which performance remains comparable or slightly lower across settings. Thus, in relation to RQ1, these results indicate that incorporating rationales tends to enhance overall predictive performance rather than only changing accuracy alone. This trend has been largely consistent across contemporary text classification literature, which is primarily English-centric Gao et al. (2024); Pruthi et al. (2022); Herrewijnen et al. (2024), and to the best of our knowledge, this study is the first to investigate it for Telugu. We refer readers to Table 1 for a consolidated overview of the models’ performance. For a detailed overview of the selection of hyperparameters for both types of training, we refer readers to Appendix K.
Model Accuracy F1 AUROC
mBERT wo/r 0.664 0.655 0.820
w/r 0.666 0.662 0.830
IndicBERT wo/r 0.713 0.711 0.879
w/r 0.715 0.718 0.886
MuRIL wo/r 0.698 0.699 0.865
w/r 0.712 0.715 0.873
XLM-R wo/r 0.705 0.705 0.864
w/r 0.692 0.695 0.867
Te-BERT wo/r 0.716 0.712 0.876
w/r 0.729 0.730 0.884
Table 1: Models’ performance for both types of training. wo/r: without rational, w/r: with rational
7.2 Aligning With Human Understanding
Given the low-resource nature of Telugu, we do not attempt mechanistic or latent-space analyses (e.g., neuron activations or representational geometry), which are difficult to interpret without strong linguistic priors and diagnostic tools. Instead, we evaluate alignment through post-hoc explanations, which provide a human-interpretable view of model behavior. Our dataset includes explicit rationales collected from multiple annotators for every instance, offering a grounded representation of human reasoning. Importantly, post-hoc explanation methods are non-trainable and applied independently of the training procedure, ensuring that comparisons between label-only and rationale-supervised models are not confounded. We use plausibility metrics DeYoung et al. (2020) as a reasonable proxy for alignment, which measure how well model-generated explanations overlap with human-provided rationales. When explanations are reasonably faithful (i.e., explanations do reflect the underlying reasoning of the model Jacovi and Goldberg (2020)), higher plausibility indicates closer agreement between what the model attends to and what humans consider relevant—making plausibility a practical, interpretable proxy for human alignment in this setting.
Across models and explanation methods, rationale supervision yields consistent gains in plausibility, but not in faithfulness. For each plausibility (A, F, I) and faithfulness (C, S, L) metric, we perform paired two-sided t-tests and Wilcoxon signed-rank tests with a significance threshold of $p<0.05$. We observe statistically significant improvements for F and I under both tests, while A shows a significant effect under Wilcoxon and a marginal trend under the t-test. In contrast, none of the faithfulness metrics are statistically significant under either test, indicating no systematic change in faithfulness due to rationale supervision. This divergence is expected. Faithfulness evaluates how accurately an explanation reflects a model’s internal reasoning process, whereas rationale supervision modifies the model itself rather than the explainer. Since post-hoc explainers are fixed, non-trainable tools Rudin (2019), there is no inherent reason to expect systematic changes in faithfulness. Plausibility, by contrast, directly measures agreement with human-identified evidence. The observed improvements, therefore, indicate that rational supervision encourages models to rely on input regions that humans themselves deem relevant for sentiment decisions, making model behavior more aligned with human understanding. This directly addresses RQ2. To the best of our knowledge, contemporary literature lacks a systematic analysis of plausibility and faithfulness that is curated under a comparable experimental setup to ours, even in English-centric domains, let alone in low-resource settings. The work closest in spirit is by Carton et al. Carton et al. (2022), which reports marginal and variable improvements using a task-specific proxy faithfulness measure that may not be directly applicable to other settings, including ours. More broadly, existing studies in this area are fragmented and narrowly focused, making it difficult to draw transferable insights for our research question Gao et al. (2024). Our detailed results are provided in Appendix H.
7.3 Bias & Fairness Analysis
We analyze how alignment-oriented training interacts with social bias along gender and religion dimensions, where males and Hindus are treated as privileged groups based on the demographic distribution of native Telugu speakers. Overall, training with rationales induces only small relative changes in fairness metrics, including polarity distribution shift ($PDS$), true positive rate difference ($TPR_{diff}$), and false positive rate difference ($FPR_{diff}$), and these changes are not consistent across models or demographic groups. MuRIL and mBERT present notable exceptions in the case of religion bias, with a substantial decrease observed for MuRIL and a moderate increase for mBERT. Detailed fairness results are reported in Appendix H.
Answering RQ3, these results indicate that while alignment-oriented design choices reliably improve human alignment, they do not induce systematic or guaranteed improvements in fairness: models that become more aligned with human reasoning are not necessarily more fair. Though there’s some occasional improvement in bias reduction in our findings, that is evidently not a clear indication that alignment-oriented supervision alone is sufficient to ensure consistent fairness improvements. This outcome is expected, as our training objective explicitly targets alignment with human understanding and does not impose fairness-specific constraints. Moreover, Telugu is an extremely low-resource language, and the pre-trained models considered here are trained on substantially less and less diverse data than those for high-resource languages. As a result, the scale and nature of social biases in Telugu models may differ from those reported in English-centric studies such as HateXplain Mathew et al. (2021).
8 Conclusion and Future Direction
In this work, we argued that progress in low-resource sentiment analysis should be measured not only by accuracy, but also by how models align with human reasoning, interpretation, and social context. To this end, we introduced TeSent, a richly annotated Telugu sentiment dataset with human rationales, and TeEEC, a controlled large corpus for systematic analysis of bias evaluation. Using these resources, we showed that incorporating human rationales consistently improves alignment with human explanations and often enhances predictive performance, and we also found that alignment-oriented training alone does not yield systematic improvements in the fairness landscape. Taken together, our resources and findings provide a comprehensive empirical foundation for studying performance, alignment, and fairness in this low-resource setup. For future work, we aim to mechanistically interpret how training with rationales shapes learning dynamics & fairness for Telugu.
Funding Information
We are thankful to all the annotators, without whose effort and feedback this work would not have been possible. We also thank the Department of Computer Science at SRM University AP for funding the project. This work was supported by the UG-PG Seed Grant Scheme of SRM University AP, India, under Grant No. SRMAP/UG-PG/SEED/2024-25/30.
Appendix
Appendix A Data Split
Figure 2 illustrates the data split used for constructing the TeSent dataset.
Figure 2: Data Split for TeSent
Appendix B Detailed Related Works Section
@@
Initial resources for sentiment classification in English language comprises of product reviews Pang et al. (2002); Maas et al. (2011), and the relevance of sentiment classification of product reviews will probably never diminish. However, extensive use of social media platforms has drawn attention to understanding sentiment polarity of short social media posts Socher et al. (2013); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021). Casual nature of such texts makes them extremely noisy containing grammatical errors, typos, an abundance of non-standard text, slang and code-mixing. This posed new challenges in curating benchmark datasets and addressing the problem itself. Mostly, sentiment classification has been considered as a binary Socher et al. (2013); Orbach et al. (2021); Pang et al. (2002); Maas et al. (2011), 3-class Dong et al. (2014), or 5-class Wang et al. (2010); Rosenthal et al. (2017) problem. While most of the popular datasets sourced the data from one particular social media platform such as Twitter Dong et al. (2014), IMDB Maas et al. (2011), tripAdvisor Wang et al. (2010), etc., datasets like Orbach et al. (2021) considered diversifying sources over multiple platforms.
@@
The lack of high-quality datasets for sentiment classification in low-resource Indic languages has significantly limited research in this area. The most commonly cited corpora include Hindi Akhtar et al. (2016), Bangla Islam et al. (2023), Marathi Kulkarni et al. (2021), multiple languages that include Telugu Doddapaneni et al. (2023) and code-mixed texts with English Patwa et al. (2020); Chakravarthi et al. (2020a, b). Mukku and Mamidi (2017) and Marreddy et al. (2022) are the only popular Telugu sentiment corpus. Among the aforementioned Indic corpora, only Islam et al. Islam et al. (2023) follow a well-defined annotation protocol, described in detail, and have curated a large-scale dataset from Bangla texts across diverse domains, collected from multiple social media platforms. The remaining studies did not diversify their data sources. A few of them have small sample sizes Mukku and Mamidi (2017); Chakravarthi et al. (2020a), some do not discuss the annotation protocol in detail Marreddy et al. (2022); Kulkarni et al. (2021); Akhtar et al. (2016), and others suffer from weak data and annotations Doddapaneni et al. (2023); Kulkarni et al. (2021); Patwa et al. (2020); Chakravarthi et al. (2020a); Mukku and Mamidi (2017). None of these works including Islam et al. Islam et al. (2023) extend their benchmarks to address explainability and fairness considerations. To the best of our knowledge, ours is the first large-scale Telugu sentiment corpus—and possibly the second among Indic languages after Islam et al. Islam et al. (2023)—comprising texts from diverse domains, sourced from multiple social media platforms, and developed using a well-defined and well-documented annotation protocol. Furthermore, it is the first Indic sentiment corpus to incorporate considerations of explainability and fairness.
@@
Explainability: Several of-the-shelf explainers have gained popularity for model auditing which are post-hoc in nature Bhatt et al. (2020). These explainers weight input tokens based on their contribution to the prediction. Since they are based on crude heuristics, these explainers are inherently imperfect Rudin (2019). Plausibility and Faithfulness are the two primary categories of evaluation measures for these explainers Lyu et al. (2024). A salient feature of our dataset is the inclusion of human-annotated rationales, which serve as ground truth for plausibility evaluation. While this has been explored to some extent in English Herrewijnen et al. (2024), our work is the first to address it in the context of an Indic language.
@@
Fairness in sentiment classification has been evaluated in English for various social biases, including gender bias Thelwall (2018); Kiritchenko and Mohammad (2018), age-related bias Díaz et al. (2018), and racial bias Kiritchenko and Mohammad (2018) among others. Attempts to evaluate fairness in Indic sentiment classification are rare, with only a few recent studies—such as those in Bangla Das et al. (2023, 2024)—curating dataset for addressing this issue. Our work is the first attempt to provide corpus and subsequently evaluate fairness in Telugu sentiment classification, and a rare effort for the same across Indic languages.
Appendix C Preprocessing Pipeline